OpenStudy (anonymous):

Simplify: (16x12) /(4x2)

OpenStudy (studygurl14):

\[\frac{16x^{12}}{4x^{2}}\] I'm guessing...

OpenStudy (studygurl14):

First do 16/4 \[\frac{4x^{12}}{x^{2}}\]

OpenStudy (studygurl14):

\[\frac{4x^{12}}{x^{2}} \rightarrow (4x^{12})\frac{1}{x^{2}}\] \[\frac{1}{x^{2}} = x^{-2}\] \[4x^{12}x^{-2}\] What next?

OpenStudy (the_fizicx99):

 Remember the exponent rules, when your dividing you subtract them.